# Generating metamers of human scene understanding

### v1

@@ 6.1 Neurally-grounded feature maps
We compared human judgments to a model whose internal representations systematically correspond to human visual processing. We employed a blur-trained AlexNet architecture (Jang and Tong, 2024), which has been specifically trained to be robust to image blur and whose internal representations exhibit strong correlations with human neural responses across visual areas from V1 to inferotemporal cortex (IT). This neurally-grounded model allowed us to isolate contributions from different stages of the visual hierarchy to metameric perception. As illustrated in Figure 4, our analysis pipeline treats early, mid, and late layers as proxies for different stages of visual processing. For each layer, we extracted feature maps from both original and generated images and computed cosine similarity to quantify alignment across the visual hierarchy. We found that as feature similarity increased at any processing level, the proportion of participants judging images as metameric also increased. This relationship held consistently across all layers of the network, from early visual features through high-level representations. The results demonstrate that metamerism spans the entire visual hierarchy rather than being confined to a single processing stage, suggesting that successful scene metamers must maintain representational alignment across multiple levels of visual processing.
Importantly, we observed distinct patterns when comparing images generated based on the random-sampling human-fixation sampling conditions. While early-layer feature similarities showed little difference between the two conditions, mid- and late-layer similarities revealed divergent trends: that human-fixation sampling maintained a consistent linear relationship between high-level feature alignment and “same” metameric judgments, while random-sampling produced an inverted-U trend for late-layer feature similarities.
@@ 6.2 Interpretable visual feature analysis
Having demonstrated that neurally grounded feature similarity aligns with human metameric judgments, we turned next to explicitly defined, interpretable visual features. To capture contributions across different levels of the visual hierarchy, we analyzed a diverse set of features: low-level (e.g., edges, Gabor filters, color), mid-level (e.g., depth cues, proto-object structure), and high-level (e.g., object, semantics, overall perceptual similarity). Because many of these features are correlated, we applied a forward stepwise regression model to identify the most predictive subset ($R^{2}=0.039$), which we focus on in the main text. Detailed contributions of each feature to the regression are provided in Appendix A.5.
Figure 5: (Left) Mid-level visual features driving metameric judgments: For metamers generated based on human-fixated locations (salmon), the preservation of monocular depth estimates in scene structure was an indicator of how more depth discrepancies yielded a decrease in metameric judgments. Additionally, when it came towards the mid-level organizational structure as seen from proto-object candidates, greater mIoU scores correlated with greater proportions of “same” metameric judgments. (Right) High-level visual features driving metameric judgments: Semantic similarity strongly predicts metameric perception, with larger DreamSim distances corresponding to reduced perceptual alignment. This result is shared with the CLIP similarity trends as well. However, these trends are less apparent when metamers were generated based on randomly-sampled locations (turquoise).
@@ 6.2.1 Low-level visual features
We compared human “same” judgments as a function of (i) Gabor filter intensities and (ii) Sobel edge density response differences between the generated and original images. By comparing normalized Gabor filter responses, of four orientations (0°, 45°, 90°, 135°), between the original and generated images, we assessed how low-level texture detection affects scene similarity judgments. Surprisingly, we found that positive differences in Gabor filter responses—where generated images showed stronger texture responses than originals—correlated with more “same” judgments. This suggests that enhanced texture definition, which makes boundaries more distinctive, increases the perceived realism of generated images, even when they differ substantially from the originals (Ho et al., 2012). We also found that greater Sobel edge density responses (Kanopoulos et al., 1988) led to greater “same” judgments, though this effect was redundant with the Gabor filter effect (see A.4).
@@ 6.2.2 Mid-level visual features
We tested two different mid-level visual features, representing local scene layout information available prior to full scene segmentation: (i) relative depth and (ii) proto-object segmentation. Depth information proved crucial for metameric perception.
We utilized the Depth Anything model (Yang et al., 2024a) to obtain depth maps from both original and generated images, then compared them using the Scale-Invariant Logarithmic (SiLog) error metric (Lee et al., 2021; Eigen et al., 2014). As discrepancies between depth maps increased, the proportion of “same” metameric judgments systematically decreased (Figure 5). This finding highlights how fundamental depth perception is to mid-level scene understanding and spatial layout representation (Verhoef et al., 2016).
We also extracted proto-object segmentations to analyze mid-level grouping structures using the conv3 layer of the blur-trained AlexNet model (Jang and Tong, 2024). These mid-layer representations are crucial for forming robust “proto-object” identities (Finkel and Sajda, 1992; Yu et al., 2014), which are initial, structured percepts that represent candidate objects by integrating visual parts and features before final semantic recognition. Greater proto-object segmentation similarity (mIoU) predicted more “same” judgments (Figure 5). This demonstrates that proto-object structures—the intermediate groupings that bridge low-level features and high-level object recognition—play a role in scene representation.
@@ 6.2.3 High-level visual features
For high-level semantic comparison, we employed both (i) CLIP (Radford et al., 2021) and (ii) DreamSim (Fu et al., 2023) as learned semantic similarity models.
For metamers generated using human-guided fixations, DreamSim served as the strongest predictor of metameric perception among all features tested. DreamSim was specifically trained on human judgments using a two-alternative forced-choice paradigm to capture human-like notions of visual similarity, with smaller DreamSim discrepancies predicting more ‘same’ responses (Figure 5, right). DreamSim’s superior explanatory power likely stems from its ability to capture mid-to-high visual features that cannot be described in language, making it a more comprehensive measure of perceptual similarity than models focused solely on semantic content.
Similarly, as semantic alignment increased between generated and original scenes, measured by CLIP representation similarity, participants were more likely to judge them as metameric (Figure 5 Right). This effect was specific to human-generated fixations and absent for randomly generated fixations, where higher CLIP similarity did not translate into more “same” metameric judgments. We suggest that this discrepancy reflects the fact that random fixations often fall on contextually irrelevant regions, exposing semantic details misaligned with participants’ internal scene representations. Together, these results indicate that metamers generated from human fixation sequences produced scenes that are better aligned with participants’ internal representations, particularly at the level of high-level semantics. Additional object-level visual feature analyses can be seen in Appendix A.4.
@@ 6.3 Foveal and peripheral features both contribute to metameric judgments
We ran an ablation experiment to isolate the contributions of foveal and peripheral conditioning in MetamerGen. We recruited 10 additional participants for a same-different task similar to the primary experiment reported above, with four second-image conditions that systematically assessed the impact of conditioning: identical original images (actual “same” images), generated images using both foveal and peripheral conditioning (as in the primary experiment), generated images using peripheral-only conditioning, and generated images using foveal-only conditioning. We found that, whereas both foveal and peripheral conditioning played a role in whether a generation becomes a metamer, the role played by peripheral conditioning was greater. As expected, the full model had the highest fool rate of 54.5%, compared to the second highest fool rate of 45.8% in the peripheral-only generation condition. Because the model learned to rely on peripheral conditioning for generating scene structure and foveal conditioning for generating the fine-grained visual information at fixated locations, images generated using only foveal conditioning tended to be easily distinguishable from original images (8.4% “same” judgments). Nevertheless, this shows that additional conditioning from foveal inputs contributes visual and semantic information that produces generations that are better aligned with human scene understanding, beyond what peripheral-only conditioning alone can achieve.
We replicated the preceding multi-level visual feature analysis under each condition of this ablation experiment, and found that (1) under equivalent levels of feature similarity, the full model is the most likely to fool participants; and (2) in general, feature similarity predicted participant judgments the best amongst full model generations, more poorly amongst peripheral-only generations, and very little amongst foveal-only generations. For more detailed results and figures, see Appendix  A.9.
@@ 7 Limitations
While MetamerGen is effective at reconstructing semantically coherent scenes from sparse visual inputs, it inherits limitations from the pre-trained Stable Diffusion model on which it is built. In our work, we identified two main limitations in the generated images: (1) difficulties producing fine-grained facial details and accurate limb articulations (Narasimhaswamy et al., 2024; Wang et al., 2025), and (2) generations of text were often unreadable (Yang et al., 2024b) even when directly fixated. To mitigate the effects of these model weaknesses on our behavioral experiment, we excluded images containing such problematic elements as inputs to MetamerGen. Including these elements would have caused participants to respond “different” due to Stable Diffusion artifacts rather than differences in their own scene representations.
@@ 8 Discussion
In this paper, we introduced MetamerGen, a latent diffusion model that generates image metamers aligned with human scene representations by combining peripheral gist with fixation-based information. While MetamerGen was trained to predict images from randomly sampled locations, we found that the scientific value of the model is maximized when conditioned on human fixations. For some applications, such as large-scale or crowdsourced experiments, random conditioning offers practical flexibility by removing the need for eye-tracking, and we found that it is indeed capable of fooling viewers. However, fixation-based conditioning better reflects human perceptual processes, reducing noise in behavioral judgments and yielding stronger correlations across all feature hierarchies (Figs. 4, 5). MetamerGen advances generative modeling by producing semantically coherent and diverse scenes from sparsely sampled inputs. We also believe it offers a powerful tool for cognitive scientists studying scene perception, enabling testing fixation-specific hypotheses on how scene representations unfold dynamically.
@@ A.10 DINOv2 versus CLIP as the vision encoder of MetamerGen
Figure 16: Image generation examples across blur levels using DINOv2 and CLIP as vision encoders. DINOv2-based peripheral generations resemble the original images more than CLIP-based generations, even at low blur levels. As the rate of downsampling decreases ($28\times 28\rightarrow 448\times 448$), DINOv2-based generations continue to show substantial improvements while CLIP-based generations exhibit minimal improvements. For the bottom two row pairs, DINOv2-based generations are able to keep the size of the plane (as well as its spatial position) intact irregardless of blur level input. However, that is not the case for the CLIP-based generations.

### v2

@@ 5.1 Neurally-grounded feature maps
Figure 4: Multi-level feature analysis pipeline using neurally-grounded model: (Top) Early, mid, and late network layers serve as proxies for different stages of processing across the hierarchy of visual brain areas. (Bottom) Results show that as feature similarity increased at these different processing levels, the proportion of participants judging generated images as metameric also increased. These effects were clearer when metamers were generated based on a viewer’s own fixated locations (salmon) than on randomly-sampled locations (turquoise).
To investigate the level of visual processing in the human brain that is most responsible for a scene becoming a metamer, we compared our behavioral judgments to an AlexNet that was trained to be robust to image blur and demonstrated to have internal representations aligned with neural responses from visual brain areas ranging from V1 to inferotemporal cortex (IT) (Jang and Tong, 2024). Using this neurally-grounded model, we isolated contributions from early to late levels of the visual hierarchy to metameric perception. As illustrated in Figure 4, our analysis pipeline treats early, middle, and late layers as proxies for different stages of visual processing. For each layer, we extracted feature maps from both original and generated images and computed cosine similarity to quantify alignment across the visual hierarchy. For attention-guided generations, we found that as feature similarity increased, the proportion of participants judging images as metameric also increased. This relationship held consistently across all layers of the network, from early visual features through high-level representations. These results suggest that metamerism spans the entire hierarchy of visual processing rather than being confined to a single processing stage, and that in order for a scene to be judged a metamer there must be broad representational alignment across low-level and high-level visual features.
Although scenes generated from actual fixations and random fixations yielded roughly the same metamer rates (29.4% and 27.7%, respectively, $p=0.24$), closer comparison hints at an interaction in the mid- and late-layer similarities. For attention-guided metamers, higher similarity to the original predicted more “same” judgments in a consistently linear trend. However, for randomly-guided generations, metamerism rates decreased when feature similarity was high. We interpret this trend as suggesting that realistic details in non-fixated regions may expose inconsistencies with the viewer’s internal scene representation that make the generated scene discriminable from the original.
@@ 5.2 Interpretable visual feature analysis
Having demonstrated that metameric judgments align with neurally grounded feature similarity, we next performed analyses aimed at making our results more interpretable. We did this by focusing on specific features likely to reflect low-, mid-, and high-level visual processing. We considered a diverse set of features: low-level (e.g., edges, Gabor filters, color), mid-level (e.g., depth cues, proto-object structure), and high-level (e.g., object, semantics, overall perceptual similarity). Because many of these features are correlated, we applied a forward stepwise regression model to identify the most predictive subset ($R^{2}=0.039$), which we focus on in the main text. Detailed contributions of each feature can be found in Appendix A.5.
Figure 5: (Left) Mid-level visual features driving metamer judgments: For metamers generated from the viewer’s own fixation locations (salmon), changes in monocular depth estimates of scene structure strongly predicted “same” judgments to generations. The alignment in proto-object segmentation between original and generated scenes, quantified by mIoU, similarly predicted metamerism rate (higher mIoU scores correlating with higher proportions of “same” judgments), although here the relationship was less pronounced. (Right) High-level visual features driving metamer judgments: Semantic similarity strongly predicts metameric scene understanding, with larger DreamSim distances corresponding to reduced perceptual alignment. CLIP similarity shows a similar trend, but not when scenes were generated from randomly-sampled locations (turquoise).
@@ 5.2.1 Low-level visual features
We compared human “same” judgments as a function of (i) Gabor filter intensities and (ii) Sobel edge density response differences between the generated and original images to assess how low-level texture features affect similarity judgments. We used normalized responses from Gabor filters oriente at 0°, 45°, 90°, and 135°. Surprisingly, we found that positive differences in Gabor filter responses—where generated images showed stronger texture responses than originals—correlated with more “same” judgments. This suggests that enhanced texture definition, which makes boundaries more distinctive, increases the perceived realism of generated images, even when they differ substantially from the originals (Ho et al., 2012). We also found that greater Sobel edge density responses (Kanopoulos et al., 1988) led to greater “same” judgments, though this effect was redundant with the Gabor filter effect (see A.4).
@@ 5.2.2 Mid-level visual features
We tested two different mid-level visual features that represent local scene layout information prior to full scene segmentation: (i) relative depth and (ii) proto-object segmentation.
We used Depth Anything (Yang et al., 2024a) to obtain depth maps from the original and generated images and compared them using Scale-Invariant Log error (SiLog) (Lee et al., 2021; Eigen et al., 2014). As shown in the leftmost plot from Figure 5, as discrepancies between depth maps increased, the proportion of “same” metameric judgments systematically decreased. This pronounced effect of depth on metamerism rate suggests that a mid-level representation of the depth relationships in a scene, essential for capturing scene layout (Verhoef et al., 2016), is an important factor in determining whether a generated scene becomes a metamer.
To analyze the effect of mid-level grouping on metamer rates, we extracted proto-object segmentations using the conv3 layer of the blur-trained AlexNet (Jang and Tong, 2024). These mid-layer representations are important to the formation of robust “proto-objects” (Finkel and Sajda, 1992; Yu et al., 2014), which are the pre-semantic building blocks of visual objects created by grouping local features into simple shapes. We found that greater mIoU in proto-object segmentation predicted more “same” judgments (Figure 5), demonstrating that similarity in proto-object segmentation also affects scene metamerism. Of the two mid-level features tested, however, the effect of depth seemed the clearer.
@@ 5.2.3 High-level visual features
To analyze the effect of high-level semantic features on metamerism rate, we used both (i) CLIP (Radford et al., 2021) and (ii) DreamSim (Fu et al., 2023) as learned semantic similarity models. Higher CLIP values indicate greater semantic alignment between original and generated scenes, whereas for DreamSim greater alignment is indicated by lower values.
DreamSim distance provides the clearest evidence for high-level semantic alignment contributing to a metameric judgment. We found that smaller distance values predict more “same” responses to generations (Figure 5, rightmost plot). This is likely because DreamSim was specifically trained on human judgments using a two-alternative forced-choice paradigm to capture human-like notions of visual similarity. CLIP also predicted this relationship between semantic alignment and the likelihood of a generation becoming a metamer (Figure 5). However, metamerism rate increased with semantic alignment only when generations were based on the viewer’s own fixations. When generations were from random fixations, higher CLIP similarity did not predict more “same” metameric judgments. We suggest that this interaction may be due to random fixations often falling on contextually irrelevant regions, thereby exposing semantic details that are not aligned with the viewer’s internal scene representation. A similar interaction between own-fixation and random conditions appears in DreamSim, but only at high distances. Together, we interpret our results to mean that scenes generated from a viewer’s attention have better high-level semantic alignment with the viewer’s internal scene representation and are therefore more likely to be confused as originals. Additional object-level visual feature analyses can be found in Appendix A.4.
@@ 5.3 Foveal and peripheral features both contribute to metameric judgments
We ran an ablation experiment to isolate the contribution of foveal and peripheral conditioning in MetamerGen. We recruited 10 additional viewers to participate in a same-different task similar to the primary experiment reported above, but with four second-image conditions that systematically assessed the impact of conditioning: identical original images (true “same” responses), generated images using both foveal and peripheral conditioning (as in the primary experiment), generated images using peripheral-only, and generated images using foveal-only conditioning. We found that while both foveal and peripheral conditioning played a role in whether a generation becomes a metamer, the role played by peripheral conditioning was greater. As expected, the full model had the highest metamer rate of 54.5%, compared to the second highest rate of 45.8% in the peripheral-only generation condition. The metamer rate in the foveal-only condition was only 8.4%. This much lower rate was due to the generations in that condition correctly capturing fixated details but differing from originals substantially in the non-fixated regions. In contrast, generations in the peripheral-only condition were more likely to be judged metamers because they better captured the global scene structure and layout. Nevertheless, this ablation analysis shows that the additional conditioning from foveal inputs, when combined with the peripheral conditioning, contributes visual and semantic information that leads togenerations that are better aligned with human scene understanding than generations produced from peripheral conditioning alone.
We replicated the multi-level visual feature analysis from Section 5.2 for each condition of this ablation experiment and found that: (1) under equivalent levels of feature similarity, generations from the full model are most likely to produce metamers, and (2) in general, feature similarity predicted participant judgments the best among full model generations, more poorly among peripheral-only generations, and very little among foveal-only generations. For more detailed results and figures, see Appendix  A.9.
@@ 6 Limitations
While MetamerGen is effective at generating semantically coherent scenes from sparse visual inputs, it inherits limitations from the pre-trained Stable Diffusion model on which it is built. In our work, we identified two main limitations: (1) difficulties producing fine-grained facial details and accurate limb articulations (Narasimhaswamy et al., 2024; Wang et al., 2025), and (2) generations of text were often unreadable (Yang et al., 2024b) even when directly fixated. To mitigate the effects of these model weaknesses on our behavioral experiment, we excluded images containing such problematic elements. Including these elements would have caused participants to respond “different” due to Stable Diffusion artifacts rather than differences in their own scene representations.
@@ 7 Discussion
This paper introduces MetamerGen, a latent diffusion model for generating images that are aligned with human scene representations enough to become scene metamers. MetamerGen demonstrates that sparse conditioning based on peripheral gist and fixation-specific details is sufficient to produce human-aligned scene representations. The work described in this paper should be of broad interest to communities ranging from cognitive scientists to machine learning researchers. For cognitive scientists, MetamerGen becomes a powerful tool for exploring the human representation of scenes. It enables the testing of fixation-specific hypotheses about what a person’s scene understanding would be given their allocation of attention during scene viewing. For machine learning researchers, MetamerGen advances generative modeling by producing from sparsely sampled inputs semantically coherent scenes that are demonstrated to align with human scene representations. This greater human alignment will contribute to next-generation assistive technologies capable of interacting more naturally with humans.
@@
Acknowledgments
We would like to thank the National Science Foundation for supporting this work through awards 2123920 and 2444540 to GJZ, and the National Institutes of Health through their award R01EY030669, also to GJZ. AL is supported by NSF-GRFP award 2234683. AG is supported by NSF grants IIS-2123920, IIS-2212046 awarded to DS. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the funding agency.
@@ A.10 DINOv2 versus CLIP as the vision encoder of MetamerGen
Figure 16: Image generation examples across blur levels using DINOv2 and CLIP as vision encoders. DINOv2-based peripheral generations resemble the original images more than CLIP-based generations, even at low blur levels. As the rate of downsampling decreases ($28\times 28\rightarrow 448\times 448$), DINOv2-based generations continue to show substantial improvements while CLIP-based generations exhibit minimal improvements. For the bottom two row pairs, DINOv2-based generations are able to keep the size of the plane (as well as its spatial position) intact irregardless of blur level input. However, that is not the case for the CLIP-based image generations.